5. Prove that if A1, A2, ... An and B are sets, then (A. – B) U (A2 – B) U... U (An – B) = (A, U A, U... U An) – B.
Answer #1

solu kon : Here to prove (A1-B) U (A2-B) D--- U CAn-B) = (D1UAZ U-- UPO) –B let 3C E (01-B) U (A2-B)U -- O(An-B) = XE (Ai-B)